# Quantum feedback induced entanglement relaxation and dynamical phase transition in monitored free fermion chains with Wannier-Stark ladder

### v2

@@ Abstract
In recent years, measurement-induced entanglement transitions (MIET) have attracted significant attention. However, the dynamical transition associated with the feedback-induced skin effect, which exhibits a wealth of intriguing phenomena, has not been fully understood. In this work, we investigate a dynamical phase transition in a tilted free-fermion chain under measurement-feedback protocols, emphasizing particle density and entanglement entropy dynamics. We reveal a feedback-induced skin effect, enhanced by the Wannier-Stark ladder potential, which creates localization at one boundary and generates an effective pseudo-edge under periodic conditions. The observables show a two-stage evolution: a rapid initial logarithmic growth followed by decay into an area-law steady state. Using finite-size scaling analysis, we pinpoint the critical transition and provide an intuitive physical picture linking it to feedback-driven suppression of quantum jump fluctuations. While the Wannier-Stark ladder delays relaxation without changing universality, strong gradients can directly induce area-law entanglement, bypassing the dynamical transition. This work provides a effective supplement to the dynamical transition proposed in Ref. [1]. It provides valuable references for linking the dynamical understanding of the role feedback plays in MIET.
@@ I Introduction
In this work, we investigate a dynamical phase transition in a tilted (Wannier-Stark laddered) free-fermion chain under monitoring-feedback protocols, with a focus on the dynamics of particle density profiles and entanglement entropy. We discover a feedback-induced skin pumping effect, where the presence of the Wannier-Stark ladder extends the evolution time required for the initial density profile to transition into the final domain-wall configuration. Under periodic boundary conditions, this potential generates an effective boundary or pseudo-edge, facilitating similar localization even in the absence of physical edges. Our analysis of entanglement entropy indicates a two-step evolution: a rapid initial increase following logarithmic scaling, subsequently transitioning into a gradual decay toward an area-law entangled steady state. Using multiple observables combined with finite-size scaling, we accurately pinpoint the critical transition separating these dynamical regimes. Furthermore, we offer an intuitive picture connecting this behavior to feedback-driven suppression of quantum jump fluctuations. Importantly, the Wannier-Stark ladder delays system relaxation but does not alter the dynamical phase transition’s universality class. Nevertheless, strong gradient potentials can independently lead the system directly into area-law entanglement, thus circumventing an explicit dynamical transition. This work effectively complement the unresolved issues regarding the mechanism and criticality of dynamical phase transitions highlighted in Ref. [1]. It provides valuable references for the possible connection between MIET and non-Hermitian dynamics.
This paper is organized as follows. Section II introduces the model along with the monitoring-feedback protocol. In Section III, we present a comprehensive analysis of the dynamical relaxation behavior in our system. We begin by examining the trajectory-averaged relaxations, focusing on the time evolution of the average density distribution. Next, we provide numerical evidence for the entanglement dynamical transitions under varying tilts, and discuss the evolution of other relevant observables during the relaxation process. Finally, using finite-size scaling, we analyze the critical parameters of this dynamical phase transition, which further confirms the universal nature of the observed behavior. Section IV provides an explanation for this dynamical phase transition, summarizes the paper, and outlines future research directions. Additional numerical calculation data and discussions are included in the Appendix.
@@ II.1 Model
We choose the Néel state as the initial state, i.e., $\left|\psi(t=0)\right\rangle=\left|0101...01\right\rangle$. Since both $\hat{H}_{\text{eff}}$ and $\hat{L}_{l}$ are quadratic, we can utilize the Gaussian state to efficiently simulate each trajectory evolution of $\left|\psi(t)\right\rangle$. We note the observables are averaged 200-800 trajectories, represented by the variable with an overline (e.g., trajectory-averaged density distribution $\overline{\langle\hat{n}_{l}\rangle}$). We describe the evolution of the observables using the rescaled time $t/\tau$, where $\tau$ is a dimensionless coefficient proportional to $L$.
@@ III.1 Trajectory-averaged relaxations
Figure 2: Panels (a)-(d) decipt the evolution of the trajectory-averaged density distribution for $L$ = $128$. (a) OBC, $\Delta$ = 00; (b) OBC, $\Delta$ = $0.6$; (c) PBC, $\Delta$ = 00; (d) PBC, $\Delta$ = $0.6$. Panels (e) and (h) show the evolution of the $\overline{S}_{L/2}$ for different $L$ under OBCs and PBCs. The green solid lines represent the absense of the Wannier-Stark ladder ($\Delta$ = 00), while the orange solid lines denote its presence ($\Delta$ = $0.6$). The colors rangeing from light to dark correspond to $L$ = $64$, $96$, $128$, $160$ and $192$. In the zoomed inset, we highlight the transition by the dashed black line. We also show the $\overline{S}_{L/2}$ versus $\log{L}$ at $t/\tau$ = $0.5$(diamond), $1.8$(square). Panels (f) and (i) illustrate the maximum bipartite entanglement entropy $\overline{S}^{\text{max}}_{L/2}$ with respect to $L$ = $64$, $96$, $128$, $160$ and $192$ (from light to dark colors). The insets show the $\overline{S}^{\text{max}}_{L/2}$ versus $\log{L}$ for different $\Delta$ under OBCs and PBCs separately. Panel (g) shows the evolution of $\overline{v}$ for different $\Delta$. Panel (j) shows the evolution of $\overline{v}$ for different $\gamma$ under $\Delta$ = 00. Both panels (g) and (h) have a system size of $64$.
@@ III.2 Entanglement relaxations and dynamical phase transitions
Now, we delve into the dynamics of the entanglement relaxation by examining the bipartite entanglement entropy, denoted as $\overline{S}_{L/2}$, across various $L$ and $\Delta$. Under OBCs, in FIG. 2 (e), $\overline{S}_{L/2}$ first sharply increases, then after the transient metastable value, slowly declines until the steady value irrespective of the presence or absence of the Wannier-Stark ladder. The entanglement relaxation with the Wannier-Stark ladder ($\Delta=0.6$) requires more time than the case without it ($\Delta=0$). Notably, there is a cross point during the decline of $\overline{S}_{L/2}$ for different $L$ and the same $\Delta$. We emphasize them with dashed lines in zoomed insets, called transition points. The transition point separates the log-law phase from the area-law phase. The inset of FIG. 2 (e) presents several examples, illustrating the relationship between $\overline{S}_{L/2}$ and $\log{L}$ when $\Delta$ is 0.0 and 0.6, respectively. In the left of the transition point the relationship between $\overline{S}_{L/2}$ and $\log{L}$ is a linear increase; but in the right, $\overline{S}_{L/2}$ remains constant regardless of $\log{L}$. For $\Delta=0$, the critical time (rescaled time $t_{c}/\tau$) is about $0.79$, which is consistent with the results in Ref. [1]; for $\Delta=0.6$, $t_{c}/\tau$approx $1.45$. As the duration of the entanglement-relaxation process increases, the transition point is delayed, corresponding to the lower zone of FIG. 1 (b) which comprises both the log-law and area-law regions. We have noticed that if $\Delta$ is significant enough and the evolution is slow enough, we suggest that there might not exist the transition point based on long-term observations. We provide examples in the Appendix C to demonstrate this viewpoint. For this case, $\overline{S}_{L/2}$ stays in the area-law phase after an initial rapid increase. This novel phenomenon is demonstrated in the upper zone of FIG. 1 (b).
@@
Figure 3: The evolution of different observables under various $\Delta$ values. The columns of the panels correspond to $\Delta=0.0$, $0.2$, $0.4$ and $0.6$, respectively. Panels (a)-(d) display the trajectory-averaged mutual information $\overline{I}_{AB}$. The selected subsystems $A$ and $B$, each with a size of $L/8$ and separated by $L/2$, are depicted in the schematic of panel (a). Panels (e)-(h) describe the evolution of the average probability $\overline{f}_{\text{skin}}$ of the evolved state being in the ideal many-body state over time. Panels (i)-(h) are about $\overline{\|D\|}_{F}$, describing the evolution of the correlations between different subsystems under the system bipartition. The curves in the figure range from light to dark, corresponding to system sizes from small to large.
@@
We investigated the time evolution of additional observables during the relaxation process, with particular emphasis on their behavior at the transition point. In FIG. 3, we display three distinct observables, with each column’s panels corresponding to the same $\Delta$ value. The black dashed line indicates the transition time, $t_{c}/\tau$, which is subsequently validated through finite-size scaling analysis. Panels (a)–(d) illustrate the evolution of $\overline{I}_{AB}$. For various system sizes $L$, we observe that $\overline{I}_{AB}$ initially increases with time, then decreases, eventually reaching zero at $t_{c}/\tau$. This indicates that once the system attains a steady state, there is no correlation between subsystems $A$ and $B$. In panels (e)–(h), we depict the evolution of $\overline{f}_{\text{skin}}$ , which quantifies the probability that the evolving state $|\psi(t)\rangle$ is in the ideal many-body state $|\psi_{\text{skin}}\rangle$. It is evident that before $t_{c}/\tau$, $\overline{f}_{\text{skin}}$ remains at 0, indicating that $|\psi(t)\rangle$ can never be in $|\psi_{\text{skin}}\rangle$. However, at $t_{c}/\tau$ this probability surges sharply, though it does not reach 1. This corresponds to the domain wall state without a distinct boundary, as depicted in FIG. 2 (a) and (b). Panels (i)–(l) depict the evolution of $\overline{\|D\|}_{F}$. This observable, which is similar in nature to $\overline{I}_{AB}$, follows an evolution pattern akin to that of $\overline{S}_{L/2}$. As shown, a crossover occurs at $t_{c}/\tau$, after which the values rapidly stabilize. Taking all of the above observables into account, we find that their critical behavior is also delayed as $\Delta$ increases.
@@ III.3 Criticalities of the dynamical phase transitions
For the free fermion model. the finite size effect is significant, and hence it is not trivial to obtain critical points which separate the log-law phase and the area-law phase in entanglement relaxation dynamics. To investigate this critical properties, we obtained scaling collapses of $\overline{S}_{L/2}$ as a function of rescaled time for $\Delta=0.0$, $0.2$, $0.4$, and $0.6$ under open boundary conditions (OBCs), all of which are shown in the lower zone of FIG. 1. In data collapse analysis, we employed the relation, $\overline{S}_{L/2}(t/\tau,L)L^{-\beta}=F[(t/\tau-t_{c}/\tau)L^{1/\nu}]$, where $F$ is a polynomial function. We identified a set of critical parameters { $t_{c}/\tau$, $\nu$, $\beta$ in Table. 1, which enable data collapse for system of different $L$, shown in FIG. 4.
Table 1: The critical information of dynamical phase transition under different $\Delta$. The initial states in the first three rows correspond to the Néel state, while the initial states in the last row correspond to the anti-Skin-like state (The state where particles accumulate at the opposite boundary).
$\Delta$ $t_{c}/\tau$ $\nu$ $\beta\times 10^{-2}$
$0.0$ $0.79(2)$ $1.23(4)$ $1.57(5)$
$0.2$ $0.89(5)$ $1.20(2)$ $2.69(3)$
$0.4$ $1.10(6)$ $1.19(4)$ $5.20(1)$
$0.6$ $1.45(4)$ $1.12(7)$ $3.12(2)$
$0.0$ $1.46(4)$ $1.23(6)$ $2.89(8)$
Figure 4: Finite-size scaling for $\overline{S}_{L/2}$ with $\overline{S}_{L/2}L^{-\beta}$ versus $(t/\tau-t_{c}/\tau)L^{1/\nu}$. Panels (a)-(d) correspond to $\Delta$ = $0.0$, $0.2$, $0.4$ and $0.6$, respectively. The system sizes $L$ = $32$, $64$, $96$, $128$ and $192$ are represented by a color gradient from light to dark, where lighter shades correspond to smaller $L$ and darker shades to larger $L$. Insets in each panel show the evolution of $\overline{S}_{L/2}$ over time for the corresponding $\Delta$. The black dashed lines in each inset represent the $t_{c}/\tau$ for the respective panel.
It is worth noting that as $\Delta$ increases, the critical time $t_{c}/\tau$ gradually increases as well. However, the critical exponents $\beta$ and $\nu$ remain largely unchanged. Therefore, we conclude that the Wannier-Stark ladder does not alter the intrinsic universality class of this dynamical phase transition. From the perspective of this critical exponent, it suggests that the system belongs to the Ising universality class.
@@ IV Discussion and Conclusion
    The resulting transition exhibits finite-size scaling in a dimensionless time variable, $t/\tau$, where $\tau\sim L$ is the characteristic transport time in a 1D open system.
@@
    A moderate Wannier-Stark ladder does not affect or alter the universality class of this dynamical phase transition.
Figure 5: Panels (a) and (e) illustrate the schematic diagrams of entanglement structures with and without feedback operations, respectively. Panels (b) and (f) describe the entanglement relaxation process for another initial state. For the initial state, the particles are mainly concentrated on the left side for panel (b) and on the right side for panel (f). The inset in (f) is finite-size scaling for $\overline{S}_{L/2}$ with $\overline{S}_{L/2}L^{-0.029}$ versus $(t/\tau-1.46)L^{1/1.23}$. Panels (c) and (g), we rescale the time and show the evolution of $\overline{S}_{L/2}$. Panels (d) and (h) illustrate the evolution of $\overline{S}_{L/2}$ for different $\gamma$. All panels (b)-(d) and (f)-(h) are in $\Delta=0$.
@@
Separately, introducing the feedback operator Eq. (3) with $\theta=\pi$ balances these two extremes. The feedback acts to erase the noise of quantum jumps, thereby pushing the system toward the area-law entangled, non-Hermitian skin-like steady state. However, the feedback effect is not instantaneous. Initially, the system remains in a metastable log-law entangled regime; only after a characteristic relaxation time does the feedback reshape the evolving state into the final skin-like state. This two-stage process is shown schematically by the solid green lines in FIG. 5 (e) and numerically in FIG 2 (e) (green lines). Early on, the system exhibits log-law entanglement, then crosses over to the area-law entangled steady state. The finite-size crossing in the time domain thus emerges naturally as a feedback-induced dynamical phase transition. Additionally, the lifetime of the metastable log-law regime is determined by how quickly the feedback pumping reshapes the system into the skin-like steady state. If the initial state is already similar to the final steady state, relaxation occurs swiftly, and the metastable window is narrow. Conversely, if the initial state is orthogonal, a longer metastable regime is observed. FIG. 5 (b) illustrates the entanglement dynamics starting from the effective Hamiltonian’s skin-like steady state: no metastable regime appears, as the system is already at equilibrium under the non-Hermitian dynamics. In contrast, FIG. 5 (f) shows the opposite case, where the initial state is a reverse skin state, fully orthogonal to the final steady state. The metastable window is significantly prolonged, and the dynamical transition time increases to $t_{c}/\tau\approx 1.46$, compared to a shorter time in the Néel-state case [FIG 2 (e), green lines]. Since the choice of the initial state does not alter the mechanism between the MIET dynamics and the non-Hermitian effective Hamiltonian dynamics, the critical exponents remain independent of the initial state. Our numerical results are shown in TABLE 1 and the inset of FIG. 5 (f).
@@
To understand the critical behavior, we note that $t/L$ (proposed in Refs. [66, 70, 1, 71]) are dimensionally confusing, as it does not carry units of time. Instead, we identify $t/\tau$ as a more natural scaling variable, where $\tau$ is the time for particles to traverse a 1D open chain of length $L$. As $\tau\sim L$, the ratio $t/\tau$ is dimensionless and suitable for finite-size scaling. When other nonlinear time rescalings are attempted [FIG. 5 (c) and (g)], signatures of the dynamical phase transition disappear, confirming that $t/\tau$ is the appropriate scaling parameter. Moreover, at higher measurement rates $\gamma$ [FIG. 5 (d) and (h)], the metastable state itself becomes area-law, rather than log-law, but we anticipate the same feedback mechanism to induce a crossover between two distinct states in the entanglement evolution. Observing this transition in practice, however, is more challenging because stronger quantum-jump noise obscures the crossover.
@@
It worth to notice that the Wannier-Stark ladder preserves the universality class of this dynamical phase transition. As discussed in Ref. [1], the localization of the effective Hamiltonian only monotonically influences the time it takes for particles to be transported to the boundary. The stronger the localization, the longer the transport time. This suggests that the criticality of this dynamical phase transition is not entirely determined by the dynamics of the effective Hamiltonian.
@@
To sum up, we systematically investigate a dynamical phase transition in monitored free fermion chain subject to feedback control. We study the evolution of the particle density distribution and entanglement entropy under different boundary conditions (open and periodic) and analyze the impact of the Wannier-Stark ladder. Firstly, we identify and characterize the feedback-induced skin effect, where the trajectory-averaged particle density becomes localized on one side of the lattice at late times. Introducing a Wannier-Stark ladder significantly prolongs the residence time in this non-equilibrium state, as the potential gradient restricts particle transport. Under periodic boundary conditions, this gradient creates a pseudo-edge, leading to a similar localization effect even without physical edges. Secondly, by analyzing the trajectory-averaged entanglement entropy, we observe a clear two-stage dynamical process: initially, entanglement rapidly grows into a metastable regime characterized by logarithmic scaling (log-law), and subsequently decays slowly into a final steady state exhibiting area-law entanglement. This transition between log-law and area-law behaviors occurs at a distinct dynamical transition point, confirmed by examining multiple observables, including subsystem mutual information, probability of the evolving state being in the ideal skin state, and subsystem correlation norms. Using finite-size scaling analysis, we precisely identify critical parameters associated with this dynamical phase transition. Importantly, we provide an intuitive and graphical explanation for the mechanism driving this dynamical phase transition: feedback gradually suppresses quantum jump noise, reshaping the system’s steady state from a measurement-induced metastable log-law entangled state into a non-Hermitian skin-like area-law state. Moreover, we discuss how the presence of the Wannier-Stark ladder affects the dynamics: while it significantly delays the relaxation process, it does not alter the universality class of the observed dynamical phase transition. Finally, we explore the fate of this dynamical transition in the presence of a strong Wannier-Stark ladder. We propose that if the gradient is sufficiently large, the system may directly relax into an area-law regime without undergoing a clear dynamical transition.

### v3

@@ Abstract
In recent years, measurement induced entanglement transitions (MIETs) have attracted significant attention. However, the dynamical transition associated with the feedback induced skin effect, which exhibits a wealth of intriguing phenomena, has not been fully understood. In this work, we investigate a dynamical phase transition in a tilted free-fermion chain under measurement-feedback protocols, emphasizing the particle density and entanglement entropy dynamics. We reveal a feedback induced skin effect, enhanced by the Wannier-Stark ladder potential, that creates localization at one boundary and generates an effective pseudo edge under periodic conditions. The observables show a two-stage evolution: a rapid initial logarithmic growth followed by decay into an area-law steady state. Using a rescaling analysis, we pinpoint the critical behavior and offer an intuitive physical picture that links it to the feedback-driven suppression of quantum jump fluctuations. The resulting entanglement dynamics appear to be governed by a system-size-dependent delay, followed by a size-independent relaxation process. This behavior is consistent with the ballistic propagation of free fermions toward a domain-wall-like steady state and does not exhibit any signatures of nontrivial criticality. This work provides an effective supplement to the dynamical transition. It provides valuable references for linking the dynamical understanding of the role feedback plays in MIETs.
@@ I Introduction
In this work, we investigate a dynamical phase transition in a tilted (Wannier-Stark laddered) free-fermion chain under monitoring-feedback protocols, with a focus on the dynamics of particle density profiles and entanglement entropy. We discover a feedback-induced skin pumping effect, where the presence of the Wannier-Stark ladder extends the evolution time required for the initial density profile to transition into the final domain-wall configuration. Under periodic boundary conditions, this potential generates an effective boundary or pseudo-edge, facilitating similar localization even in the absence of physical edges. Our analysis of entanglement entropy indicates a two-step evolution: a rapid initial increase following logarithmic scaling, subsequently transitioning into a gradual decay toward an area-law entangled steady state. Using multiple observables combined with data collapses, we accurately pinpoint the critical transition separating these dynamical regimes. Furthermore, we offer an intuitive picture connecting this behavior to feedback-driven suppression of quantum jump fluctuations. Importantly, the Wannier-Stark ladder delays system relaxation but does not alter the dynamical phase transition’s critical behavior. Nevertheless, strong gradient potentials can independently lead the system directly into area-law entanglement, thus circumventing an explicit dynamical transition. This work effectively complement the unresolved issues regarding the mechanism and criticality of dynamical phase transitions highlighted in Ref. [70]. It provides valuable references for the possible connection between MIET and non-Hermitian dynamics.
This paper is organized as follows. Section II introduces the model along with the monitoring-feedback protocol. In Section III, we present a comprehensive analysis of the dynamical relaxation behavior in our system. We begin by examining the trajectory-averaged relaxations, focusing on the time evolution of the average density distribution. Next, we provide numerical evidence for the entanglement dynamical transitions under varying tilts, and discuss the evolution of other relevant observables during the relaxation process. Finally, using rescaling analysis, we analyze the critical points of this dynamical phase transition, which further confirms the universal nature of the observed behavior. Section IV provides an explanation for this dynamical phase transition, summarizes the paper, and outlines future research directions. Additional numerical calculation data and discussions are included in the Appendix.
@@ II.1 Model
We choose the Néel state as the initial state, i.e., $\left|\psi(t=0)\right\rangle=\left|0101...01\right\rangle$. Since both $\hat{H}_{\text{eff}}$ and $\hat{L}_{l}$ are quadratic, we can utilize the Gaussian state to efficiently simulate each trajectory evolution of $\left|\psi(t)\right\rangle$. We note the observables are averaged 200-800 trajectories, represented by the variable with an overline (e.g., trajectory-averaged density distribution $\overline{\langle\hat{n}_{l}\rangle}$). We describe the evolution of the observables using the rescaled time $t/\tau$, where $\tau$ is a time scale proportional to $L$.
@@ III.1 Trajectory-averaged relaxations
Figure 2: Panels (a)-(d) depict the evolution of the trajectory-averaged density distribution for $L$ = $128$. (a) OBC, $\Delta$ = 00; (b) OBC, $\Delta$ = $0.6$; (c) PBC, $\Delta$ = 00; (d) PBC, $\Delta$ = $0.6$. Panels (e) and (h) show the evolution of the $\overline{S}_{L/2}$ for different $L$ under OBCs and PBCs. The green solid lines represent the absense of the Wannier-Stark ladder ($\Delta$ = 00), while the orange solid lines denote its presence ($\Delta$ = $0.6$). The colors ranging from light to dark correspond to $L$ = $64$, $96$, $128$, $160$ and $192$. In the zoomed inset, we highlight the transition by the dashed black line. We also show the $\overline{S}_{L/2}$ versus $\log{L}$ at $t/\tau$ = $0.5$(diamond), $1.8$(square). Panels (f) and (i) illustrate the maximum bipartite entanglement entropy $\overline{S}^{\text{max}}_{L/2}$ with respect to $L$ = $64$, $96$, $128$, $160$ and $192$ (from light to dark colors). The insets show the $\overline{S}^{\text{max}}_{L/2}$ versus $\log{L}$ for different $\Delta$ under OBCs and PBCs separately. Panel (g) shows the evolution of $\overline{v}$ for different $\Delta$. Panel (j) shows the evolution of $\overline{v}$ for different $\gamma$ under $\Delta$ = 00. Both panels (g) and (h) have a system size of $64$.
@@ III.2 Entanglement relaxations and dynamical phase transitions
Now, we delve into the dynamics of the entanglement relaxation by examining the bipartite entanglement entropy, denoted as $\overline{S}_{L/2}$, across various $L$ and $\Delta$. Under OBCs, in FIG. 2 (e), $\overline{S}_{L/2}$ first sharply increases, then after the transient metastable value, slowly declines until the steady value irrespective of the presence or absence of the Wannier-Stark ladder. The entanglement relaxation with the Wannier-Stark ladder ($\Delta=0.6$) requires more time than the case without it ($\Delta=0$). Notably, there is a crossing point during the decline of $\overline{S}_{L/2}$ for different $L$ and the same $\Delta$. We emphasize them with dashed lines in zoomed insets, called transition points. The transition point separates the log-law phase from the area-law phase. The inset of FIG. 2 (e) presents several examples, illustrating the relationship between $\overline{S}_{L/2}$ and $\log{L}$ when $\Delta$ is 0.0 and 0.6, respectively. In the left of the transition point the relationship between $\overline{S}_{L/2}$ and $\log{L}$ is a linear increase; but in the right, $\overline{S}_{L/2}$ remains constant regardless of $\log{L}$. For $\Delta=0$, the critical time (rescaled time $t_{c}/\tau$) is about $0.79$, which is consistent with the results in Ref. [70]; for $\Delta=0.6$, $t_{c}/\tau$approx $1.45$. As the duration of the entanglement-relaxation process increases, the transition point is delayed, corresponding to the lower zone of FIG. 1 (b) which comprises both the log-law and area-law regions. Our long-term observations suggest that, when $\Delta$ is sufficiently large and the evolution sufficiently slow, a transition point may not exist. We provide examples in the Appendix C to demonstrate this viewpoint. For this case, $\overline{S}_{L/2}$ stays in the area-law phase after an initial rapid increase. This novel phenomenon is demonstrated in the upper zone of FIG. 1 (b).
@@
Figure 3: The evolution of different observables under various $\Delta$ values. The columns of the panels correspond to $\Delta=0.0$, $0.2$, $0.4$ and $0.6$, respectively. Panels (a)-(d) display the trajectory-averaged mutual information $\overline{I}_{AB}$. The selected subsystems $A$ and $B$, each with a size of $L/8$ and separated by $L/2$, are depicted in the schematic of panel (a). Panels (e)-(h) describe the evolution of the average probability $\overline{f}_{\text{skin}}$ of the evolved state being in the ideal many-body state over time. Panels (i)-(l) are about $\overline{\|D\|}_{F}$, describing the evolution of the correlations between different subsystems under the system bipartition. The curves in the figure range from light to dark, corresponding to system sizes from small to large.
@@
We investigated the time evolution of additional observables during the relaxation process, with particular emphasis on their behavior at the transition point. In FIG. 3, we display three distinct observables, with each column’s panels corresponding to the same $\Delta$ value. The black dashed line indicates the transition time, $t_{c}/\tau$, which is subsequently validated through the data collapse. Panels (a)–(d) illustrate the evolution of $\overline{I}_{AB}$. For various system sizes $L$, we observe that $\overline{I}_{AB}$ initially increases with time, then decreases, eventually reaching zero at $t_{c}/\tau$. This indicates that once the system attains a steady state, there is no correlation between subsystems $A$ and $B$. In panels (e)–(h), we depict the evolution of $\overline{f}_{\text{skin}}$ , which quantifies the probability that the evolving state $|\psi(t)\rangle$ is in the ideal many-body state $|\psi_{\text{skin}}\rangle$. It is evident that before $t_{c}/\tau$, $\overline{f}_{\text{skin}}$ remains at 0, indicating that $|\psi(t)\rangle$ can never be in $|\psi_{\text{skin}}\rangle$. However, at $t_{c}/\tau$ this probability surges sharply, though it does not reach 1. This corresponds to the domain wall state without a distinct boundary, as depicted in FIG. 2 (a) and (b). Panels (i)–(l) depict the evolution of $\overline{\|D\|}_{F}$. This observable, which is similar in nature to $\overline{I}_{AB}$, follows an evolution pattern akin to that of $\overline{S}_{L/2}$. As shown, a crossover occurs at $t_{c}/\tau$, after which the values rapidly stabilize. Taking all of the above observables into account, we find that their critical behavior is also delayed as $\Delta$ increases.
@@ III.3 Criticalities of the dynamical phase transitions
For the free fermion model under open boundary conditions (OBCs), we find that the entanglement relaxation dynamics exhibits a pronounced dependence on system size. In particular, the characteristic time for entanglement growth increases approximately linearly with system size $L$, and the overall time evolution appears to follow a universal form when plotted against the rescaled time $t-(t_{c}/\tau)L$, as shown in Fig. 4. More importantly, as $\Delta$ increases, the critical time $t_{c}/\tau$ gradually increases as well. From the panel (a) to (d), corresponding $t_{c}/\tau$ values are approximately $0.79$, $0.89$, $1.10$, and $1.45$, respectively.
This form of data collapse indicates that the entanglement dynamics is governed primarily by a system-size-dependent delay, followed by a size-independent relaxation process. Such behavior is consistent with the propagation of free fermions toward a domain-wall-like steady state, and does not exhibit any signatures of non-trivial criticality. Additionally, the presence of a Stark potential of moderate strength only prolongs the time required to reach the domain-wall-like steady state.
Figure 4: Data collapses for $\overline{S}_{L/2}$ with $\overline{S}_{L/2}$ versus $t-(t_{c}/\tau)L$. Panels (a)–(d) correspond to $\Delta$ = $0.0$, $0.2$, $0.4$ and $0.6$, while the corresponding $t_{c}/\tau$ values are approximately $0.79$, $0.89$, $1.10$, and $1.45$, respectively. The system sizes $L$ = $64$, $96$, $128$, $160$ and $192$ are represented by a color gradient from light to dark, where lighter shades correspond to smaller $L$ and darker shades to larger $L$. Insets in each panel show the evolution of $\overline{S}_{L/2}$ over time for the corresponding $\Delta$. The black dashed lines in each inset represent the $t_{c}/\tau$ for the respective panel.
@@ IV Discussion and Conclusion
    The resulting transition exhibits rescaling in a dimensionless time variable, $t/\tau$, where $\tau\sim L$ is the characteristic transport time in a 1D open system.
@@
    A moderate Wannier-Stark ladder does not affect or alter the critical behavior of this dynamical phase transition.
Figure 5: Panels (a) and (e) illustrate the schematic diagrams of entanglement structures with and without feedback operations, respectively. Panels (b) and (f) describe the entanglement relaxation process for another initial state. For the initial state, the particles are mainly concentrated on the left side for panel (b) and on the right side for panel (f). The inset in (f) is the data collapse for $\overline{S}_{L/2}$ with $\overline{S}_{L/2}$ versus $t-1.46L$. Panels (c) and (g), we rescale the time and show the evolution of $\overline{S}_{L/2}$. Panels (d) and (h) illustrate the evolution of $\overline{S}_{L/2}$ for different $\gamma$. All panels (b)-(d) and (f)-(h) are in $\Delta=0$.
@@
Separately, introducing the feedback operator Eq. (3) with $\theta=\pi$ balances these two extremes. The feedback acts to erase the noise of quantum jumps, thereby pushing the system toward the area-law entangled, non-Hermitian skin-like steady state. However, the feedback effect is not instantaneous. Initially, the system remains in a metastable log-law entangled regime; only after a characteristic relaxation time does the feedback reshape the evolving state into the final skin-like state. This two-stage process is shown schematically by the solid green lines in FIG. 5 (e) and numerically in FIG. 2 (e) (green lines). Early on, the system exhibits log-law entanglement, then crosses over to the area-law entangled steady state. The finite-size crossing in the time domain thus emerges naturally as a feedback-induced dynamical phase transition. Additionally, the lifetime of the metastable log-law regime is determined by how quickly the feedback pumping reshapes the system into the skin-like steady state. If the initial state is already similar to the final steady state, relaxation occurs swiftly, and the metastable window is narrow. Conversely, if the initial state is orthogonal, a longer metastable regime is observed. FIG. 5 (b) illustrates the entanglement dynamics starting from the effective Hamiltonian’s skin-like steady state: no metastable regime appears, as the system is already at equilibrium under the non-Hermitian dynamics. In contrast, FIG. 5 (f) shows the opposite case, where the initial state is a reverse skin state, fully orthogonal to the final steady state. The metastable window is significantly prolonged, and the dynamical transition time increases to $t_{c}/\tau\approx 1.46$, compared to a shorter time in the Néel-state case [FIG. 2 (e), green lines].
@@
To understand the critical behavior, we note that $t/L$ (proposed in Refs. [65, 69, 70, 71]) are dimensionally confusing, as it does not carry units of time. Instead, we identify $t/\tau$ as a more natural scaling variable, where $\tau$ is the time for particles to traverse a 1D open chain of length $L$. As $\tau\sim L$, the ratio $t/\tau$ is dimensionless and suitable for the data collapse. When other nonlinear time rescalings are attempted [FIG. 5 (c) and (g)], signatures of the dynamical phase transition disappear, confirming that $t/\tau$ is the appropriate scaling parameter. Moreover, at higher measurement rates $\gamma$ [FIG. 5 (d) and (h)], the metastable state itself becomes area-law, rather than log-law, but we anticipate the same feedback mechanism to induce a crossover between two distinct states in the entanglement evolution. Observing this transition in practice, however, is more challenging because stronger quantum-jump noise obscures the crossover.
@@
It worth to notice that the Wannier-Stark ladder preserves the critical behavior of this dynamical phase transition. As discussed in Ref. [70], the localization of the effective Hamiltonian only monotonically influences the time it takes for particles to be transported to the boundary. The stronger the localization, the longer the transport time. This suggests that the criticality of this dynamical phase transition is not entirely determined by the dynamics of the effective Hamiltonian.
@@
To sum up, we systematically investigate a dynamical phase transition in monitored free fermion chain subject to feedback control. We study the evolution of the particle density distribution and entanglement entropy under different boundary conditions (open and periodic) and analyze the impact of the Wannier-Stark ladder. Firstly, we identify and characterize the feedback-induced skin effect, where the trajectory-averaged particle density becomes localized on one side of the lattice at late times. Introducing a Wannier-Stark ladder significantly prolongs the residence time in this non-equilibrium state, as the potential gradient restricts particle transport. Under periodic boundary conditions, this gradient creates a pseudo-edge, leading to a similar localization effect even without physical edges. Secondly, by analyzing the trajectory-averaged entanglement entropy, we observe a clear two-stage dynamical process: initially, entanglement rapidly grows into a metastable regime characterized by logarithmic scaling (log-law), and subsequently decays slowly into a final steady state exhibiting area-law entanglement. This transition between log-law and area-law behaviors occurs at a distinct dynamical transition point, confirmed by examining multiple observables, including subsystem mutual information, probability of the evolving state being in the ideal skin state, and subsystem correlation norms. Importantly, we provide an intuitive and graphical explanation for the mechanism driving this dynamical phase transition: feedback gradually suppresses quantum jump noise, reshaping the system’s steady state from a measurement-induced metastable log-law entangled state into a non-Hermitian skin-like area-law state. Moreover, we discuss how the presence of the Wannier-Stark ladder affects the dynamics: while it significantly delays the relaxation process, it does not alter the critical behavior of the observed dynamical phase transition. Finally, we explore the fate of this dynamical transition in the presence of a strong Wannier-Stark ladder. We propose that if the gradient is sufficiently large, the system may directly relax into an area-law regime without undergoing a clear dynamical transition.